Expedition 306 officially began with the first line ashore on Ponta Delgada, Azores (Portugal), at 1805 h on 2 March 2005. The third Ponta Delgada port call in a row for the JOIDES Resolution, which included refueling and restocking bentonite and attapulgite bulk drilling mud, was concluded at 0806 h on 9 March, ~1.8 days behind schedule because of problems with the passive heave compensator (PHC) seal replacement and severe weather that significantly hampered (and slowed) the PHC repair work. Jeff Fox, Director, Science Services, IODP-Texas A&M University (TAMU), attended the port call accompanied by the new TAMU Dean of Geosciences, Dr. Bjorn Kjerfve. We completed the transit from Ponta Delgada to Site U1312 averaging 10.3 kt over the 344 nmi distance.Site U1312
We arrived at Site U1312 (proposed Site IRD4A) at 1730 h on 10 March and spudded Hole U1312A at 0830 h on 11 March. Core 1H was recovered in full (10.08 m), suggesting a seafloor depth of 3533.0 mbsf (Table T2). Because of excessive heave (5 m), initial coring conditions were not optimal and a more realistic seafloor depth estimate was obtained later in Hole U1312B. APC coring, utilizing nonmagnetic core barrels, continued to a depth of 237.5 mbsf. The first several cores (1H through 10H to 95.0 mbsf) had questionable shear pressures, and Core 1H required two wireline runs to recover because of a sheared overshot pin. The swell height decreased after Core 10H, and coring system performance improved accordingly. Coring in Hole U1312A was terminated after recovering Core 25H. The drillover technique was required for recovery of Cores 23H through 25H, and all core barrels fully stroked. Drillover for the last core required 2 h in the semi-indurated white ooze, and Core 23H was recovered bent. Coring may have continued further; however, the risk to the equipment was significant, the time required for further advancement was reaching a diminishing return, and the co-chief scientists were concerned about getting Hole U1312B initiated while heave conditions remained low (~1.5 m). In Hole U1312A, we cored 237.5 m, recovering 248.07 m (104.45%).Hole U1312B
After the seafloor was cleared, the vessel was offset 25 m to the northwest of Hole U1312. Hole U1312B was spudded at 2115 h on 12 March, with the bit positioned at a depth of 3528.0 mbsf, 5.0 m higher than at Hole U1312A. Core barrel 1H advanced 9.5 m and recovered 3.92 m of sediment, placing the seafloor depth for Hole U1312B at 3533.6 mbsf. After successfully achieving a good mudline core, piston coring advanced to 231.9 mbsf. The core line failed at the rope socket while attempting to recover Core 18H, necessitating a wireline fishing trip for the core barrel and sinker bar assembly. Coring continued through Core 25H (231.9 mbsf) as weather conditions progressively deteriorated, leading to another damaged core line at the rope socket and twisted piston rods. Before repairs were completed, weather conditions deteriorated to a point that precluded additional coring. A wind shift and increased velocity forced a heading change to maintain position over the hole. This caused excessive roll as the ship was exposed to multiple large swells coming from different directions. Because the electrical supervisor expressed concern about the ability to keep the ship on location, the drill string was pulled out of the hole, clearing the seafloor at 0245 h on 14 March. The ship was allowed to drift off location to minimize vessel motion during the remainder of the pipe trip. By 1700 h on 14 March, the ship was secured and began the slow dynamic positioning move into the prevailing seas back over the drill site. At 0430 h on 15 March positioning beacon SN 2199 (15 kHz, 211 dB) was recovered, officially ending operations for Hole U1312B and Site U1312. In Hole U1312B, we cored 231.9 m, recovering 236.84 m (102.08%).Transit to Site U1313
We departed the Azores Islands for Site U1313 (proposed Site IRD3A) at 0645 h on Saturday, 26 March, after spending 8 days in the lee of the islands from severe weather conditions. The 277 nmi transit to Site U1313 began with sea-state conditions still somewhat marginal but expected to continuously improve over time. The transit was completed at an average speed of 8.9 kt, and we arrived at Site U1313 at 1425 h on Easter Sunday, 27 March.Site U1313
Hole U1313A was spudded at 0005 h on 28 March, establishing a rig floor-corrected seafloor depth of 3423.3 meters below rig floor (mbrf). Core 1H recovered 5.2 m of sediment, indicating a seafloor depth of 3412.3 mbsl. APC coring, utilizing nonmagnetic core barrels, continued through Core 31H. Alloy steel barrels were utilized for Cores 32H and 33H, and drillover of stuck APC barrels was required for these last two cores. Tensor core orientation was used for all cores beginning with Core 3H from a depth of 14.7 mbsf. All core barrels fully stroked except for the last one, Core 33H. The cored interval for this hole was 308.6 m, and 319.64 m of sediment was recovered (103.6%). The drill string was pulled clear of the seafloor at 1300 h on 29 March, concluding operations in Hole U1313A.Hole U1313B
Hole U1313B was offset 25 m due north (000°) from Hole U1313A and was spudded at 1430 h on 29 March. The bit was positioned at a depth of 3421.0 mbrf, or 2.0 m deeper than Hole U1313A, and Core 1H recovered 5.90 m of sediment, establishing a rig floor-corrected seafloor depth of 3424.6 mbrf. APC coring continued through Core 20H to a depth of 186.4 mbsf. At that time, operations personnel were advised that the core breaks between Holes U1313A and U1313B were possibly becoming aligned. As a precaution to ensure adequate overlap between holes, the driller was advised to advance the bit 2.0 m further before shooting Core 21H. Coring then continued until 1930 h on 30 March, through Core 32H to a total depth of 302.4 mbsf. Tensor core orientation was used for all cores beginning with Core 3H from a depth of 15.4 mbsf. All core barrels fully stroked, and no drillover was required in this hole. Nonmagnetic core barrels were used in the recovery of all 32 cores. The cored interval for this hole was 300.4 m, and 306.54 m was recovered (102.0%).
After completing coring operations, Hole U1313B was prepared for logging. A wiper trip was conducted from total depth to 74.9 mbrf and back to total depth without difficulty. No fill, drag, or overpull was detected. The hole was swept with one final 30 bbl mud sweep, and the lockable float valve (LFV) go-devil was pumped down to lock open the LFV for the wireline logging tools. The hole was displaced with 140 bbl of sepiolite mud, and the drill pipe was retracted to 80.7 mbsf. The triple combo tool string was deployed (including the GPIT and MGT tools) to within 2.0 m of the bottom of the hole. We successfully logged 220 m of the open hole and then the interval within the pipe to the seafloor. The scheduled deployment of the FMS-sonic tool string was cancelled because the LDEO Logging Staff Scientist felt that the resistivity signal was too weak to make the deployment worthwhile. After recovery of the wireline tools, new software for the Schlumberger wireline heave-compensated logging winch was briefly tested. The software was tested for ~45 min under a variety of conditions, and the results appeared to be "very promising." The Schlumberger logging sheaves were rigged down, and the drill string was pulled clear of the seafloor at 1310 h on 31 March, ending Hole U1313B.Hole U1313C
The drillship was offset 25 m due east (90°) of Hole U1313B. The bit was positioned at a depth of 3417.0 mbrf, or 2.0 m shallower than for Hole U1313A. Hole U1313C was spudded at 1440 h on 31 March. Core 1H recovered 2.71 m and established a rig floor-corrected seafloor depth of 3423.8 mbrf. APC coring with nonmagnetic core barrels continued through Core 30H, and alloy steel barrels were used for the last two cores (31H and 32H). Tensor core orientation was used for all cores beginning with Core 4H from a depth of 21.7 mbsf. Drillover was required for recovery of Cores 29H through 32H. All core barrels fully stroked except for the last three (Cores 30H through 32H). The bit was advanced by 8.73 m prior to shooting Core 32H. The drill string was pulled clear of the seafloor at 2140 h on 1 April, officially ending Hole U1313C. The cored interval for this hole was 293.4 m, and 305.79 m of sediment was recovered (104.2%).Hole U1313D
The drillship was offset 25 m due south (180°) of Hole U1313C. The bit was positioned at a depth of 3423.0 mbrf, or 4.0 m deeper than for Hole U1313A. Hole U1313D was spudded at 2255 h on 1 April. Core 1H was recovered in full (9.80 m); therefore, the seafloor depth of 3423.0 mbrf was considered suspect. A valid seafloor depth was not critical for this hole because three good seafloor measurements had already been obtained on the three previous holes and the emphasis was to maximize recovery in Core 1H. APC coring continued without incident in Hole U1313D through Core 16H to a total depth of 152.0 mbsf. Tensor core orientation was used for all cores, beginning with Core 3H from a depth of 19.0 mbsf. All core barrels fully stroked, and no drillover was required. Nonmagnetic core barrels were used for recovery of all 16 cores. The cored interval for this hole was 152.0 m, and 159.27 m of sediment was recovered (104.8%).
The drill string was pulled clear of the seafloor at 1310 h on 2 April, and, during the pipe trip positioning, beacon SN 2199 was recovered at 1430 h. After we recovered and stored the drill string, the ship was secured for transit, all thrusters and hydrophones were pulled, and at 2030 h, the ship was switched from dynamic positioning to cruise mode and got under way for alternate Site U1314 (proposed Site GAR1B).Transit to Site U1314
We arrived at Site U1314 (proposed Site GAR1B) early Thursday morning, 7 April, after a 943 nmi transit from Site U1313. Positioning beacon SN 2199 (15.0 kHz, 211 dB) was deployed at 0330 h, and operations at Site U1314 officially began.Site U1314
Hole U1314A was spudded at 0950 h on 7 April. Core 1H recovered 1.95 m of sediment and established a rig floor-corrected seafloor depth of 2810.6 mbrf. APC coring continued without incident in Hole U1314A until Core 28H at a depth of 258.4 mbsf. The overshot pin sheared during the attempt to recover Core 28H. Unfortunately, the driller did not initially realize that the core barrel was not being recovered and proceeded to drill downhole in preparation for the next core. After recovering the core barrel on the second wireline run, the APC system was found to be heavily damaged, including overtorqued piston rod connections and other damage. Because it was going to take an hour or more to repair and rebuild the APC system, the co-chief scientists elected to terminate coring in Hole U1314A and proceed with the next hole. After repairing the APC, the drill string was pulled clear of the seafloor at 1100 h on 8 April, officially ending Hole U1314A.
Tensor core orientation was used for all cores beginning with Core 4H from a depth of 20.9 mbsf. Drillover of stuck APC barrels was not required, and all core barrels fully stroked. Nonmagnetic core barrels were used to recover all 28 cores. Problems with swollen or imploded core liners were more prevalent in this hole than experienced previously at the other sites. Of 28 cores recovered, 4 liners suffered some swelling (Cores 12H, 15H, 16H, and 23H) and 6 core liners either shattered (Core 20H) or imploded (Cores 18H, 22H, 26H, 27H, and 28H). In addition, four barrels mechanically sheared during deployment (Cores 2H, 4H, 6H, and 11H). These problems were most likely heave induced and led to more disturbed and lower-quality core than normally experienced with the APC system. The cored interval for this hole was 258.4 m, and 267.34 m of sediment was recovered (103.5%).Hole U1314B
The drillship was offset 25 m due north (000°) of Hole U1314A. The bit was positioned at a depth of 2806.0 mbrf, or 3.0 m deeper than Hole U1313A, to offset the core breaks between cores. Hole U1314B was spudded at 1220 h on 8 April, and Core 1H recovered 4.02 m and established a rig floor-corrected seafloor depth of 2811.5 mbrf. APC coring with nonmagnetic core barrels continued without incident in Hole U1314B through Core 30H to a total depth of 279.5 mbsf. Tensor core orientation was used for all cores beginning with Core 4H from a depth of 23.0 mbsf. All core barrels fully stroked, and no drillover was required in this hole. The cored interval for this hole was 279.5 m, and 285.36 m of sediment was recovered (102.1%). The drill string was pulled clear of the seafloor at 1155 h on 9 April, officially ending Hole U1314B. Although there still were some core liner incidents at this hole, core quality was improved. Of the 30 cores recovered, 4 liners were recovered swollen (Cores 15H, 21H, 22H, and 26H) and another 3 were imploded at the bottom end (Cores 9H, 19H, and 27H). The overshot pin sheared once on Core 6H, requiring a second wireline run to recover the core barrel.Hole U1314C
The drillship was offset 25 m due east (90°) of Hole U1314B. The bit was positioned at a depth of 2809.0 mbrf, or 6.0 m deeper than Hole U1314A, again to offset the core breaks between cores. Hole U1314C was spudded at 1320 h on 9 April, and Core 1H recovered 8.22 m and established a rig floor-corrected seafloor depth of 2810.3 mbrf. APC coring with nonmagnetic core barrels continued in Hole U1314C through Core 22H to a total depth of 207.7 mbsf. During this period, the weather deteriorated as a major gale approached our location and brought high winds and steadily building seas and swells with it. By Sunday morning, we had long-period heave cycles in excess of 5 m and storm-force winds were forecast for the Site U1314 location by late Sunday or early Monday. Tensor core orientation was used for all cores beginning with Core 3H from a depth of 17.7 mbsf. All core barrels fully stroked, and no drillover was required in this hole. The cored interval for this hole was 207.7 m, and 212.93 m of sediment was recovered (102.5%). The drill string was pulled clear of the seafloor at 0915 h on 10 April, and positioning beacon SN 2199 was recovered during the pipe trip, officially ending Site U1314 at 1030 h. After we recovered and stored the drill string, the ship was secured for transit and all thrusters and hydrophones were pulled. At 1545 h on 10 April, the ship was switched from dynamic positioning to cruise mode and began the sea voyage to Site U1315 (Site 642 in the Vøring Plateau).Transit to Site U1315
We arrived at Site U1315 (proposed Site 642) on the morning of 15 April after a 1190 nmi transit from Site U1314. By 1120 h, the positioning beacon was deployed and operations at the site officially began.Site U1315
After we verified the seafloor depth with the television/sonar system (1283.0 mbrf), the drill string was recovered back to the drillship and we prepared to assemble and deploy the "elevated" reentry cone assembly and 103/4 inch casing string. Following assembly, the drill string with the reentry cone structure, 103/4 inch casing, and the mud motor and underreamer drilling assembly were tripped to the seafloor, spudding Hole U1315A at 2130 h on 15 April. The base of the reentry cone reached the seafloor at 0515 h in the morning on 16 April, resulting in an average rate of penetration of 21.9 m/h. The drill string was released from the reentry cone/casing assembly at 0550 h and then recovered back to the ship.
The drill string was reassembled with a cementing bottom-hole assembly and reentry cleanout bit and tripped to the seafloor, and Hole U1315A was reentered at 2026 h on 17 April. A 5 bbl, 15.8 ppg cement plug was displaced to bottom. The pipe was pulled clear of the seafloor/reentry cone, and the vessel was offset 30 m south where the drill pipe was thoroughly circulated clean. After the cement set, we reentered the hole and lowered the pipe, tagging the top of the cement at ~164.2 mbsf or ~11.6 m above the casing shoe, which was within 1.6 m of our 10.0 m target height for the top of the cement column. After the casing string was displaced with bentonite mud, the drill string was retrieved in preparation for the CORK deployment in Hole U1315A.CORK Deployment
The CORK running tool was made up to a 2.0 m drill collar pup joint and was laid out on the rig floor. The antitorsion ring, a recent modification that is designed to prevent accidental release of the tool during deployment, was installed along with the seal stinger. The Hole U1315A CORK head was moved from the core tech shop roof to the starboard side of the pipe stabber using the number 2 crane. From there, the load was transferred to two tugger lines and the head was placed through the rotary table with the bushings removed. The head was supported using 133/8 inch casing slips with a dog collar installed above. The running tool was picked up with the drawworks elevators and was easily installed on the CORK head.
Two stands of 81/4 inch drill collars were made up to the top of the running tool, and the CORK assembly was lowered to the seafloor. During the pipe trip, the subsea television/vibration-isolated television (VIT) frame was being lowered when the video was suddenly lost. The VIT was recovered, and troubleshooting identified a cleared power circuit breaker resulting from water leakage. In addition, a broken strand of outer armor was discovered ~360 m above the cable head termination. The resulting snarl in the armor was cut off, and the ends were secured. Once the cable head was on deck, we discovered that water had entered the oil-filled section of the cable head. The cable head connection was repaired rather than spending costly time completely reheading the cable. The VIT was ready to deploy once again at 0345 h on 19 April. A total of 3.25 h was required to make the necessary repairs.
Hole 1315A was reentered for the third and final time at 0448 h on 19 April. Space out was tight given the short length of the CORK stinger (20.85 m). For thermistor string deployment, the CORK head was left 9.0 m shy of landing out in the 16 inch casing hanger. This left ~12.3 m of stinger in the hole and allowed the pipe to be hung off at the rotary table to break the drill pipe connection and deploy the thermistor string.Thermistor String Deployment
A total of 5.75 h was required to make up and deploy the thermistor string. This string was specially designed for long-term monitoring of the upper 150 m of the sediment column from the seafloor down. The thermistor string was premade up with a 1/4 inch diameter Spectra (Kevlar) rope attached to the thermistor cable with tie wraps and duct tape. The Spectra line (153.84 m) was designed to carry the load as a tension member and was terminated with a ~250 lb sinker bar (3.73 m). The thermistor cable was plugged into a battery pack/data logger assembly (0.97 m) at the top. This package was suspended from another short section of Spectra rope (2.22 m) with thimbles at each end and the XN latch assembly (0.31 m to landing shoulder). Cable "grips" were used to lift the thermistor string in ~5060 ft lifts using a double sheave assembly and two tugger lines. This was the same process used on earlier thermistor deployments; however, this time, there were problems. For reasons yet to be fully understood, at the conclusion of each lift, the thermistor cable tensed up and carried the load and the Spectra line would go slack. On most lifts, the load would eventually transfer back to the Spectra line as designed; however, on the last two lifts, the load never did transfer. Fortunately the Hole U1315A thermistor string was relatively short, and the suspended load was fairly light. To rectify the problem, we cut loose the remaining tie wraps and duct tape. We then retied and retaped the cable and rope together as the remainder of the string was deployed. At the end, where the rope was to be attached to the bottom of the data logger/battery assembly, we had a surplus of 9 ft of Spectra rope that had to be cut off, leaving the final rope length at 151.10 m. Once made up, the thermistor string assembly was deployed via wireline at 1030 h on 19 April. The XN latch assembly was landed without incident. The latch was jarred down for setting, and a 4000 lb overpull was taken to verify proper latch-in. The sinker bars were recovered via wireline, and by 1215 h we were ready to land and release the CORK assembly.CORK Landing and Release
The drill string was lowered the remaining 9.0 m (~1 knobby joint), and the CORK head landed out at exactly the correct pipe depth. About 5000 lb were put down, and then ~8000 lb of overpull was taken to verify that the CORK head was latched. This weight was slacked off, and a small amount of right hand-torque was applied to the string. The CORK running tool was visually observed to rotate slightly, and when the string was picked up, the running tool lifted cleanly off the head. Installation of the Hole U1315A CORK was officially completed as of 1237 h on 19 April.
The top drive was racked back and the drill string was recovered, clearing the rig floor at 1605 h. During the pipe trip, positioning beacon SN 2039 was recovered by 1440 h. The upper guide horn or "piccolo" was then reinstalled, ending operations for Hole U1315A.Site 642
Search for Hole 642E Reentry Cone
During the trip out of the hole, the drillship moved back to the original Site 642 prospectus coordinates. From there, a search began for the Hole 642E reentry cone. A 100 x 100 m box pattern search in 15 m swaths was initiated. The operations report for ODP Leg 104 indicated that the reentry cone base was left ~1.0 m below the seafloor, leaving ~1.52 m of cone above. Other entries in the report indicated that the last few reentries were difficult and the crew questioned whether future reentries would be possible because they felt that the cone may have been buried in drilled cuttings. It should be noted, however, that all reentries at that time were made using sonar and were not aided by any subsea television capability. The search for the reentry cone began at 2045 h on 19 April and extended through the night without success. The following morning, the ship returned to a target that was not close to where the reentry cone should have been; however, an obvious man-made object was visible on the seafloor and was detected on sonar as well. The object appeared to be an old-style reentry cone reflector. Four of these reflectors were mounted on the older-style reentry cones in the past. Convinced that the cone was buried, we spent several hours attempting to define where a cone might be relative to the single reflector that was visible. Ultimately a stab into the seabed was made with the drill string; however, this was to no avail as the driller noted drill string resistance after lowering the drill string only 9 m into the seabed. Upon reflection, we decided that the object was not the object of our search; the sonar should have identified the remaining cone and reflectors even if submerged. Considering this, coupled with the fact that we were not even close to the area that the cone should have been, we decided to continue on with the search. In so doing, we revisited all that we knew about the location of the target cone and realized that we had started our search pattern at a longitude of 2°55.7'E, failing to recognize that Hole 642E had in fact been spudded at 2°55.8'E, a full tenth of a minute off. After entering new offsets into the dynamic positioning system, we finally located the reentry cone on sonar and then ultimately with the subsea television system as well. The cone was fully visible and was not submerged or covered in cuttings. The rim and all four reflectors could be clearly identified, and the cone rim appeared to be, as reported, 1.5 to 2.0 m above the seafloor. Using the new undithered Global Positioning System capabilities of the ship, the ultimate coordinates for Hole 642E are 067°13.1850'N latitude by 002°55.7789'E longitude. The cone was actually located 548 m south of Holes 642A and 642B on a bearing of 173°. According to the documentation, the cone should have been located 450 m to the southeast (a bearing of 135°). The drill string was spaced out, and Hole 642E was reentered at 1305 h on 20 April, a total of 16.5 h after the search was initiated.Wireline Logging
The primary goal of our return to Hole 642E was to obtain a high-resolution continuous temperature log of the hole that had been left undisturbed for nearly 20 y. This hole was drilled to a total depth of 1229.4 mbsf during Leg 104. It was left cased (with 113/4 inch 54 lb/ft casing) 62.5 m into basement, placing the casing shoe at 371.5 mbsf. To minimize disturbance in the hole, the end of the pipe was placed at only 15.3 mbsf. The LDEO TAP tool was deployed with the standard triple combo (Dual Induction Tool model E [DITE]/Hostile Environment Litho-Density Sonde [HLDS]/Accelerator Porosity Sonde [APS]/Hostile Environment Gamma Ray Sonde [HNGS]) tool string at 1710 h. A good temperature log was obtained on the first run downhole; however, ledges and/or chunks of basalt falling in from above the tools proved to be problematic. After the tools were "mouse trapped" between two such zones for nearly an hour, they were recovered, having only reached a depth of ~1888 mbrf (~599 mbsf) or ~228 m into the open hole below the casing shoe. The first tool suite was recovered at 2225 h on 20 April.
After rigging down the first tool suite, the second suite of tools consisting of the FMS-sonic tool string were made up and deployed. These tools were run in at 0135 h on 21 April and were recovered at 0700 h after reaching a depth of ~1878 mbrf (~589 mbsf) or ~218 m into the open hole.
The FMS-sonic tool string was rigged down, and a final run was made only to 1200 mbsf to test the MGT pressure case. This tool leaked during a logging run earlier in the expedition; however, nothing unusual was found to explain the leak. After cleaning the pressure case thoroughly, new O-rings were installed and the body was deployed without electronics to test its pressure integrity. The tools were run in at 0750 h and recovered by 0900 h. The limited test (relatively shallow depth) was successful, and the MGT was recovered without any signs of leakage.
All wireline tools and the Schlumberger logging sheaves were rigged down by 1000 h, and preparations began for drill string recovery. The bit cleared the seafloor/reentry cone at 1005 h, and by 1310 h, all drill pipe had been recovered and all drill collars were laid out to the main deck tubular rack. During the pipe trip, positioning beacon SN 2199 was recovered on deck at 1100 h. The rig floor was secured, all thrusters and hydrophones were raised, and the drillship was under way for Dublin, Ireland, by 1315 h on 21 April 2005.
Next Section | Table of Contents